OpenStudy (anonymous):

how do i find the exact solutions to 2sin^2(X)+3sin(x)-2=0

OpenStudy (solomonzelman):

let sin(x)=a and then solve the quadratic. need more help ?

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

what about the sin^2(x) what is that

OpenStudy (jdoe0001):

\(\bf 2{\color{blue}{ sin}}^2(x)+3{\color{blue}{ sin}}(x)-2=0\qquad 2{\color{blue}{ a}}^2+3{\color{blue}{ a}}-2=0\) notice, it's just a quadratic, so you can just factor it, or use the quadratic formula

OpenStudy (solomonzelman):

sin²x becomes a²

OpenStudy (solomonzelman):

(I would just do it, without let this equal that )

OpenStudy (abmon98):

(2sinx-1)(sinx+2)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

what about 2sin^2(x)cos(x)=cos(x) on the interval of 0,2pie